# Is Translation All You Need? A Study on Solving Multilingual Tasks with Large Language Models

### v1

@@ 3 Translation for Real User Queries
The results of NLP tasks demonstrate that translating into English is a straightforward yet efficient method to address tasks in non-English languages. However, these benchmarks typically focus on specific linguistic aspects, which may not fully encapsulate the breadth and complexity of real-world user queries that often span an extensive range of topics and require nuanced comprehension. Moreover, these benchmarks are often constructed by translating the original English data into other languages Shi et al. (2022); Ponti et al. (2020); Conneau et al. (2018); Yang et al. (2019); Hasan et al. (2021). This approach leads to datasets that are not truly challenging, as they miss the rich culture-specific elements crucial for truly nuanced language understanding for different languages.
To assess the impact of translation on real-world queries, we extract user requests from ShareGPT datasets. We select 10 languages ranging from high to low resources and randomly sample 50 requests for each language from the first round of the dataset. To evaluate the quality of the responses, we use GPT-4-Turbo (gpt-4-turbo-0125) as the judge. The prompt for the judge is shown in Figure 6 in the Appendix, which is adapted from Zheng et al. (2023). With this prompt, each response will get a score from 1 to 10. Since the queries can be in various formats, we only compare two promoting strategies: 1) original queries; and 2) translated queries with Google Translate API. For the second option, we will translate the output back to the original language for consistency. We select the two best models on NLP tasks for evaluation: ChatGPT and Llama-2-70B-Chat.
3.1 Results and Discussions
In our comparison, we analyzed the scores of two sets of responses generated by the same model, calculating the win rate for each language. The results are shown in Figure 3 and we make the following observations: 1) ChatGPT shows different behaviours for high-resource languages and low-resource languages. For Japanese, Chinese, and Spanish, prompting with original queries has a higher win rate. For lower-resource languages, translation is often a better option. 2) For Llama-2-70B-Chat, translation has a higher win rate for all the languages. This echoes the fact that Llama-2-70B-Chat is an English-centric LLM. Even though translation may lose some information, the gain of better understanding after translation dominates the performance.
The distinct behaviours exhibited by Llama-2-70B-Chat and ChatGPT underscore the inherent differences between the LLMs. Llama-2-70B-Chat is predominantly English-centric, as evidenced by its improved performance in both NLP tasks and real user queries when input is translated into English. Conversely, ChatGPT demonstrates characteristics of a “multilingual LLM”, as shown in Figure 1(b), showcasing its broader linguistic capacity. Nevertheless, ChatGPT still relies on translation to bolster its performance for lower-resource languages, indicating room for improvement in its approach to truly multilingual processing.
To ascertain the extent to which answering user queries necessitates an understanding of local cultural nuances, we employed GPT-4-Turbo, utilizing a specially crafted prompt to analyze user queries across multiple languages, as shown in Figure 7 in the Appendix. The results, illustrated in Figure 4, reveal that a significant portion of the queries—ranging from 18% to 50% for each language—require knowledge of local culture for accurate responses. This variation underscores the rich cultural elements embedded within the data, affirming our hypothesis that queries containing culture-specific information pose a greater challenge for the LLMs.
@@ 3.1 Results and Discussions
3.2 Analysis
Figure 4: The percentage of the questions that necessitate local cultural knowledge.
ChatGPT has shown promising performance in some high-resource languages. We extend the experiment to investigate the effectiveness of multilingual LLMs in handling culture-specific tasks, aiming to assess how well these models grasp cultural nuances and idiomatic language. Through our experiment, we seek insights into the adaptability of LLMs across diverse cultural contexts.
3.2.1 Experiment Setup
We select the following models to represent various categories of LLMs and assess the potential differences in their response to translation:
  • ChatGPT and llama-2-70B-chat, which are the strongest models in our previous experiments.
  • Qwen1.5-72B-Chat Bai et al. (2023) and Yi-34B-Chat [^4] https://github.com/01-ai/Yi. These two open-source models demonstrate strong English and Chinese capabilities. We aim to test if they excel more in Chinese than in translating Chinese into English.
Please note that except ChatGPT, the results for the other three models are obtained by calling API from Together [^5] https://api.together.xyz/. We chose open-source models which are larger than 30B since we hypothesise that larger models contain more knowledge.
For the evaluation dataset, we choose M3Exam Zhang et al. (2023a), as the questions require strong multilingual proficiency and cultural knowledge to perform well. For the subject categories, we choose language and social science, which are more likely to contain language-dependent knowledge. For each language, we select up to 500 samples for evaluation. For the prompting strategies, we follow (Zhang et al., 2023a) and use Monolingual, EN-Instruct and EN-Translation, which correspond to Native-Basic, EN-Basic  and Trans-Google in Section 2, for comparison.
3.2.2 Results and Discussions
@@
Based on the results shown in Figure 5, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance for all languages. This observation aligns with the conclusions drawn in the study by Zhang et al., 2023a. The effectiveness of translation largely depends on whether translation errors outweigh any potential gains in better comprehension. 2) Translation helps Llama-2-70B-chat in all the languages, suggesting that the model’s underperformance is due to poor language understanding rather than limitations of cultural knowledge. 3) Qwen1.5-72B-Chat and Yi-34B-Chat excel in Chinese proficiency. The translation process hurts Chinese performance, highlighting the significant influence of translation errors on comprehension. Despite this, translation may still enhance performance in other languages, notably for Yi-34B-Chat.
From this experiment, it is evident that while translation serves as a straightforward and initially effective strategy to enhance the performance of English-centric LLMs on non-English tasks, it falls short of being optimal. Directly training LLMs in specific languages can yield superior results by eliminating the potential for translation errors.
4 General Discussions
In this work, we categorize multilingual tasks into two groups: language-independent tasks and language-dependent tasks. Among the NLP benchmarks discussed in Section 2, such as MGSM, XCOPA, XNLI, PAWS-X, and MKQA, are parallel benchmarks. This means that the examples maintain semantic equivalence across all languages, categorizing them as language-independent tasks. Language-independent tasks are typically easier to evaluate across different linguistic groups, as they rely on universal cognitive abilities, like reasoning and comprehension, rather than specific linguistic knowledge. In contrast, tasks involving real user queries and culture-related questions, discussed in Section 3, necessitate models to possess cultural and contextual understanding, categorizing them as language-dependent tasks. By distinguishing between language-dependent and language-independent tasks, we aim to promote fair and accurate evaluations across diverse linguistic contexts, enhancing the development and assessment of NLP systems in multilingual environments.
Corresponding to the two types of tasks, multilingual LLMs must develop dual competencies:
General Language Skills:
Essential for understanding and engaging with both task types, these skills enable LLMs to process and comprehend content across multiple languages, facilitating basic instruction interpretation and response generation.
Specific Language Skills:
These skills are crucial for tasks demanding an understanding of cultural nuances and linguistic subtleties, such as interpreting culturally specific queries.
By equipping LLMs with these comprehensive skills, we ensure their effectiveness in a wide array of multilingual contexts, enhancing both linguistic adaptability and cultural sensitivity.
The benefits of translation on NLP tasks demonstrate the dominance of English-centric LLMs. Even top models like ChatGPT, perform better through translation. In handling user queries, ChatGPT excels in native languages compared to translation, particularly in high-resource languages but struggles with low-resource ones. When testing on the M3Exam questions, Qwen1.5-72B-Chat and Yi-34B-Chat perform quite well in the native language for Chinese questions but have no such advantages for other languages. This dependency on translation highlights the challenges in developing LLMs that can equally handle a wide range of languages, especially those with limited resources. Such findings prompt further research and development efforts aimed at enhancing the multilingual efficacy of LLMs, striving for models that can understand and generate text across languages with minimal reliance on translation.
@@ 5 Conclusion
In this work, we perform a comprehensive assessment of LLMs across diverse multilingual tasks. These tasks include traditional NLP challenges, real user queries, and culture-specific tasks. We explore the use of translation-based methods as a strategy to overcome the limitations inherent in English-centric LLMs, thereby facilitating enhanced comprehension. However, we find that for tasks heavily reliant on language-specific nuances, employing prompts in the native language emerges as a more effective approach. This method is particularly adept at capturing the subtleties and intricacies unique to each language. Our findings suggest that the development of robust multilingual LLMs presents a more effective solution for addressing the challenges of multilingual tasks, especially for language-dependent tasks. Therefore, we advocate for the prioritization of research and development efforts towards the creation of strong multilingual large language models.
6 Limitations
@@ Acknowledgements
This research is supported, in part, by DAMO Academy through DAMO Academy Research Intern Program and Alibaba-NTU Singapore Joint Research Institute (JRI), Nanyang Technological University, Singapore. Chaoqun Liu extend his gratitude to Interdisciplinary Graduate Programme and School of Computer Science and Engineering, Nanyang Technological University, Singapore, for their support.
Appendix A Additional Experiments
A.1 NLP Tasks
Here are the detailed descriptions of the NLP tasks:
@@ A.1 NLP Tasks
The prompts of EN-Basic for each task are shown in Table 3, which are adapted from Huang et al., 2023. The prompt templates for other prompting strategies and the instructions for output formats are designed according to the descriptions in Section 2.1.3.
A.2 Additional Results
@@
XCOPA 500
Here is a premise: {premise}. What is the {question}? Help me pick the more plausible option: -choice1: {choice1}, -choice2: {choice2}
XNLI 500
{premise} Based on previous passage, is it true that {hypothesis}? 1: Yes, 2: No, or 3: Maybe?
PAWS-X 500
Sentence 1: {sentence1} Sentence 2: {sentence2} Question: Does Sentence 1 paraphrase Sentence 2? 1: Yes, 2: No?
MKQA 500
Answer the question in one or a few words in {target_language}: {question}?
XL-Sum 500
{article} Summarize the article.
@@
Table 4: Average scores of the high-resource languages and low-resource languages for the six benchmarks in zero-shot setting.
@@
Figure 6: The LLM-as-a-judge prompt for GPT-4-Turbo.
Figure 7: Prompt to check whether answering a request needs local cultural knowledge.

### v2

@@ 3 Translation for Real User Queries
The results of NLP tasks demonstrate that translating into English is a straightforward yet efficient method to address tasks in non-English languages. However, these benchmarks typically focus on specific linguistic aspects, which may not fully encapsulate the breadth and complexity of real-world user queries that often span an extensive range of topics and require nuanced comprehension. Moreover, these benchmarks are often constructed by translating the original English data into other languages Shi et al. (2022); Ponti et al. (2020); Conneau et al. (2018); Yang et al. (2019); Hasan et al. (2021). This approach leads to datasets that are not truly challenging, as they miss the rich culture-specific elements crucial for truly nuanced language understanding for different languages. To assess the impact of translation on real-world queries, we extract user requests from ShareGPT [^3] https://sharegpt.com/, a website to share real conversations with ChatGPT.
3.1 Experiment Setup
We selected 10 languages, ranging from high to low resource, and randomly sampled 100 requests for each language. However, for Romanian (ro), Ukrainian (uk), and Norwegian (no), we sampled 53, 98, and 53 requests respectively, due to the limited number of samples available from the source dataset. Since the queries can be in various formats, we only compare two promoting strategies: 1) original queries; and 2) translated queries with Google Translate API. For the second option, we translate the output back to the original language for consistency. To evaluate the quality of the responses, we use GPT-4o [^4] https://openai.com/index/hello-gpt-4o/(gpt-4o-2024-05-13) as the judge. The prompt for the judge is shown in Figure 8 in the Appendix, which is adapted from Zheng et al. (2023). With this prompt, each response will get a score from 1 to 10.
@@ 3.1 Experiment Setup
3.2 Main Results
We compared the scores of two response sets from the same model, calculating the win rate for each language. The results are shown in Figure 5, leading to the following observations: 1) ChatGPT’s performance varies across languages. For high-resource languages like Japanese, Chinese, and Spanish, original queries have a higher win rate. In contrast, for low-resource languages, the effectiveness of translation can be either better or worse, depending on the specific languages involved. 2) For Llama-2-70B-Chat, translation has a higher win rate for all languages, reflecting its English-centric nature. Despite potential information loss, the improved understanding after translation still enhances performance.
Llama-2-70B-Chat and ChatGPT exhibit distinct behaviors, reflecting their inherent differences. Llama-2-70B-Chat, being English-centric, performs better with translated inputs. Conversely, ChatGPT shows certain characteristics of a “Multilingual LLM”, as shown in Figure 1(b), mainly for high-resource languages, indicating the potential for improvement in true multilingual processing.
To determine if answering user queries requires local cultural knowledge, we used GPT-4o with a specially crafted prompt to analyze queries in multiple languages (Figure 9 in the Appendix). Results in Table 14 in the Appendix show that 30% to 74% of queries per language require cultural knowledge, highlighting the rich cultural elements in the data. Further analysis of the ShareGPT subsets requiring local cultural knowledge is in Appendix A.2.
3.3 Analysis and Discussions
Based on the previous results, ChatGPT and Llama-2-70B-chat both tend to be English-centric but ChatGPT demonstrates certain behaviors of a "Multilingual LLM". An ideal "Multilingual LLM" should exhibit equivalent performance across all languages. Consequently, we broaden our analysis to include non-English-centric LLMs and assess their performance across various tasks.
How do non-English-centric LLMs perform on culture-related tasks?
To investigate the behaviors of different LLMs on culture-related tasks, we select another two LLMs: Qwen1.5-72B-Chat Bai et al. (2023) and Yi-34B-Chat AI et al. (2024), which are not English-centric. These two open-source models demonstrate strong capabilities in both English and Chinese. Therefore, we can check whether they demonstrate multilingual behaviors in Chinese, as illustrated in Figure 1(b).
For the evaluation dataset, we choose M3Exam Zhang et al. (2023a), as the questions are real-world natural data from different languages instead of translating from English data and require strong multilingual proficiency and cultural knowledge to perform well. For example, the question about a Swahili proverb shown in Figure 2(b) requires local knowledge to answer correctly. We select the language and social science subject categories, which likely contain more native cultural knowledge, and evaluate up to 500 samples per language. For the prompting strategies, we follow (Zhang et al., 2023a) and use Monolingual, EN-Instruct and EN-Translation, which correspond to Native-Basic, EN-Basic  and Trans-Google in Section 2, for comparison.
@@ 3.3 Analysis and Discussions
Based on the results shown in Figure 6, we have the following observations: 1) For ChatGPT, translation may not always result in improved performance. This observation aligns with the conclusions drawn in the study by Zhang et al., 2023a. The effectiveness of translation largely depends on whether translation errors outweigh any potential gains in better comprehension. 2) Translation helps Llama-2-70B-chat in all the languages, suggesting that the model’s underperformance is due to poor language understanding rather than limitations of cultural knowledge. 3) Qwen1.5-72B-Chat and Yi-34B-Chat excel in Chinese proficiency. The translation process hurts Chinese performance, highlighting the significant influence of translationese on comprehension. Despite this, translation may still enhance performance in other languages, notably for Yi-34B-Chat, indicating that they are far from ideal multilingual LLMs.
How do non-English-centric LLMs perform on NLP tasks?
Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
Qwen1.5-72B-Chat
Native-Basic 78.8 93.0 55.8 71.8 36.6 41.3 62.9
EN-Basic 77.2 97.0 73.0 73.0 32.7 39.7 65.4
Native-CoT 83.2 95.8 46.4 72.2 35.8 39.5 62.1
EN-CoT 81.6 97.2 71.2 70.6 34.9 38.6 65.7
XLT 78.4 97.8 77.4 67.6 20.8 35.3 62.9
Trans-Google 81.6 94.6 63.8 68.4 45.7 31.3 64.2
Trans-NLLB 58.8 88.2 61.4 70.4 32.0 28.5 56.5
Yi-34B-Chat
Native-Basic 63.2 92.6 46.0 43.6 13.4 36.9 49.3
EN-Basic 66.8 93.6 52.6 74.6 15.5 35.1 56.4
Native-CoT 65.2 91.8 42.6 43.6 13.0 36.6 48.8
EN-CoT 70.0 93.6 48.2 74.8 12.1 33.1 55.3
XLT 56.0 93.2 69.2 65.6 7.5 31.3 53.8
Trans-Google 71.2 94.0 49.6 70.8 24.5 36.3 57.7
Trans-NLLB 56.0 86.6 48.8 68.2 22.9 28.5 51.8
Table 3: Scores of the two non-English-centric LLMs on NLP tasks for the Chinese language. The best result for each model is in bold.
As shown in Figure 2(b), for an ideal multilingual LLM, prompting in native languages should still have advantages over translation if the tasks are less dependent on languages. To test the hypothesis, we evaluate Qwen1.5-72B-Chat and Yi-34B-Chat on the NLP tasks as discussed in Section 2.1.1. We only evaluate them in Chinese since the two models are optimized for this language.
The results are displayed in Table 3. Trans-Google remains competitive among various prompting strategies, achieving the best average scores for Yi-34B-Chat, which surpasses our expectations. The possible reason could be that while both models are optimized for Chinese, their performance in Chinese still lags behind their proficiency in English. Nevertheless, We have the following special observations for the two models. 1) For Qwen1.5-72B-Chat, the best strategy is EN-CoT instead of Trans-Google. We hypothesize that this prompting strategy utilizes the model’s bilingual abilities and simultaneously avoids translationese. 2) Both LLMs perform better with Native-Basic for the XL-Sum dataset. We hypothesize that the dataset is more language-dependent than other tasks as it is created by considering the local context instead of simply translating from the English version Hasan et al. (2021). 3) The translation benefits are less pronounced than those of ChatGPT and Llama-2-70B-Chat. For example, the gap between Trans-Google and Native-Basic on MGSM(Chinese) for the two models are 2.8% and 8%. The values for ChatGPT and Llama-2-70b-Chat are 37.2% and 16%, respectively, which are significantly larger.
How do different LLMs handle multilingual prompts?
To further understand the differences between English-centric LLMs and non-English-centric LLMs, we analyze the layerwise language distribution for Llama-2-7B-Chat and Qwen1.5-7B-Chat, using the method proposed by Zhao et al., 2024. We decode the embedding after each layer and identify each token into different languages with CLD3555https://github.com/google/cld3. As shown in Figure 7, the two LLMs process Chinese prompts differently. While the hidden representations of Qwen1.5-7B-Chat are mainly in Chinese, those of Llama-2-7B-Chat are in various other languages. We hypothesize that processing the information in native without conversion avoids the information loss, making it more suitable for processing multilingual tasks.
(a) Llama-2-7B-chat
(b) Qwen1.5-7B-Chat
Figure 7: Layerwise language distribution for (a) Llama-2-7b-chat and (b) Qwen1.5-7B-Chat with Chinese prompts.
4 Related Work
Multilingual Evaluation.
Since the release of ChatGPT, the evaluation of LLMs has attracted the attention of the research communityQin et al. (2023a); Bang et al. (2023). Shi et al., 2022 evaluated LLMs on MGSM and found that the models demonstrated strong multilingual reasoning capabilities, even for low-resource languages. Bang et al., 2023 evaluated ChatGPT on 23 datasets covering 8 NLP tasks. They found that ChatGPT failed to generalize its capabilities to non-Latin scripts. To cover tasks, Ahuja et al., 2023 evaluated ChatGPT and GPT-4 on 16 NLP datasets across 70 languages and compared them with state-of-the-art non-autoregressive models. Concurrently, Lai et al., 2023 evaluated ChatGPT on 7 different tasks across 37 diverse languages. However, these evaluations are primarily limited to standard NLP tasks and largely overlook real-world scenarios and cultural knowledge Fung et al. (2024), which are crucial for understanding the practical applicability of LLMs.
Multilingual Prompting Strategies.
The translate-test is a popular technique used to refine the performance of multilingual NLP benchmarks Conneau et al. (2018); Ponti et al. (2020); Artetxe et al. (2023); Moghe et al. (2023). In the era of LLMs, various strategies have been developed to enhance the performance of LLMs using multilingual datasets. Shi et al., 2022 discovered that EN-CoT outperforms Native-CoT. Huang et al., 2023 introduced cross-lingual-thought prompting to minimize language disparities. In parallel, Qin et al., 2023b introduced cross-lingual prompting, and Etxaniz et al., 2023 suggested self-translate to elevate their performances. Effective in translating prompts into English, these methods excel in NLP tasks but remain uncertain in real-world applications. Their success hinges on the English-centric nature of the LLMs. Our study evaluates translation effectiveness across NLP tasks, real user queries, and non-English-centric LLMs, revealing the limitations of these methods.
@@ 5 Conclusion
We have conducted a thorough evaluation of LLMs in various multilingual tasks. These tasks include traditional NLP benchmarks, real user queries, and culture-related tasks. Even though translation-based methods are simple and effective strategies to overcome the limitations inherent in English-centric LLMs, they are not optimal for all scenarios, highlighting the necessity of more comprehensive multilingual evaluation. The experiment on non-English-centric LLMs and culture-related tasks demonstrates that employing prompts in the native language emerges as a more effective approach. This method is particularly adept at capturing the subtleties and intricacies unique to each language. The challenge of the setting is that it requires LLMs to be proficient in various languages, calling for the prioritization of research and development efforts toward the creation of strong multilingual LLMs.
Limitations
@@ Acknowledgements
This research is supported, in part, by DAMO Academy through DAMO Academy Research Intern Program and Alibaba-NTU Singapore Joint Research Institute (JRI), Nanyang Technological University, Singapore. Chaoqun Liu extend his gratitude to Interdisciplinary Graduate Programme and College of Computing and Data Science, Nanyang Technological University, Singapore, for their support.
Appendix A Appendix
A.1 Translation for NLP Tasks
This section presents more details about the setups and results for the experiments on NLP tasks.
A.1.1 Details about NLP Benchmarks
Here are the detailed descriptions of the NLP benchmarks:
@@ A.1.1 Details about NLP Benchmarks
A.1.2 More LLMs for Experiment
Besides ChatGPT and Llama-2-70B-Chat, we have also evaluated the NLP tasks with the following models:
  • Llama-2-13B-chat, which is a chat model in Llama-2 family Touvron et al. (2023).
  • bloomz-7b1, which is a model fine-tuned with multiple tasks, including some multilingual tasks Muennighoff et al. (2023).
A.1.3 More Details about Prompt Strategies
An example of various prompting strategies is shown in Table A.1.4. The prompts of EN-Basic for each task are shown in Table 5, which are adapted from Huang et al., 2023. The translation template for self-translate with LLMs is:
Translate the following question from {language} to English:
{question}
Don’t answer the question, just translate it!
The prompt templates for other prompting strategies and the instructions for output formats are designed according to the descriptions in Section 2.1.3.
A.1.4 Additional Results
@@
Original Question {CJK*}UTF8gbsn制作一件袍子需要 2 匹蓝色纤维布料和这个数量一半的白色纤维布料。它一共需要用掉多少匹布料
Native-Basic {Original Question }
{CJK*}UTF8gbsn您的最终答案的格式应为:"答案: <阿拉伯数字>".
EN-Basic {Original Question }
You should format your final answer as "Answer: <Arabic numeral>".
Native-CoT {Original Question }
{CJK*}UTF8gbsn让我们一步步思考。
{CJK*}UTF8gbsn您的最终答案的格式应为:"答案: <阿拉伯数字>".
EN-CoT {Original Question }
Let’s think step by step in English.
You should format your final answer as "Answer: <Arabic numeral>".
XLT I want you to act as an arithmetic reasoning expert for Chinese.
Request: {Original Question}
You should retell the request in English.
You should do step-by-step answer to obtain a number answer.
You should step-by-step answer the request.
You should tell me the answer in this format ’Answer :’.
Trans-Google Crafting a robe requires 2 bolts of blue fiber cloth and half that amount of white fiber cloth. How many pieces of fabric will it use in total?
Let’s think step by step.
You should format your final answer as "Answer: <Arabic numeral>".
Trans-NLLB To make a robe, two pieces of blue fiber and half of that amount of white fiber are needed. How many pieces of fabric does it take to make?
Let’s think step by step.
You should format your final answer as "Answer: <Arabic numeral>".
@@
XCOPA 500 Here is a premise: {premise}. What is the {question}? Help me pick the more plausible option: -choice1: {choice1}, -choice2: {choice2}
XNLI 500 {premise} Based on previous passage, is it true that {hypothesis}? 1: Yes, 2: No, or 3: Maybe?
PAWS-X 500 Sentence 1: {sentence1} Sentence 2: {sentence2} Question: Does Sentence 1 paraphrase Sentence 2? 1: Yes, 2: No?
MKQA 500 Answer the question in one or a few words in {target_language}: {question}?
XL-Sum 500 {article} Summarize the article.
@@
Table 6: Average scores of the high-resource languages and low-resource languages for the six benchmarks in zero-shot setting. The results of PAWS-X and XL-Sum for bloomz-7b1 are not considered since it was already pre-trained on these tasks. The best result for each model is in bold.
@@
Model Prompt type de ru fr zh es ja sw th bn te avg
ChatGPT Trans-ChatGPT 77.6 75.2 78.4 76.0 78.8 69.6 75.2 62.4 65.6 42.8 70.2
Llama-2-70B-Chat Trans-Llama 53.6 52.0 55.2 46.8 54.4 44.0 8.8 11.2 15.2 4.8 34.6
Table 8: Accuracy scores across various languages on the MGSM benchmark with self-translate approach.
@@
A.2 Translation for Real User Queries
The prompt used to assess the response quality is shown in Figure 8. When GPT-4o is prompted with this, it assigns a score ranging from 1 to 10 to each response. Figure 9 illustrates the prompt used to determine if responding to a request requires local cultural knowledge. The Chinese case shows that GPT-4o can identify if queries require knowledge of local culture with explanations and the final answer.
We also analyzed the performance of shareGPT subsets with cultural knowledge only. As shown in Figure 10, the behaviors across languages and models are inconsistent. ChatGPT shows different behaviors for high-resource and low-resource languages. For high-resource languages like Japanese, Chinese, and Spanish, prompting with original queries has a higher win rate. For low-resource languages, translation is often a better option. In contrast, Llama-2-70B-Chat shows a higher win rate for all languages.
Figure 8: The LLM-as-a-judge prompt for GPT-4o.
Figure 9: Prompt template to check whether answering a request needs local cultural knowledge (upper) and one Chinese example (lower).
Language ja zh es fr vi id ko ro uk no
Ratio (%) 59 58 38 41 67 55 55 74 30 57
Table 14: The percentage of the questions that necessitate local cultural knowledge.
(a) Win rate with ChatGPT (w/ cultural knowledge)
(b) Win rate with Llama-2-70B-Chat (w/ cultural knowledge)
Figure 10: Win rate comparison for each language using ChatGPT and Llama-2-70B-Chat for the subsets of shareGPT with cultural knowledge.